. The potential field 𝑉 = 2π‘₯

2𝑦𝑧 βˆ’ 𝑦

3

𝑧 exists in a dielectric medium having πœ€ = 2πœ€0 (a) Does V satisfy

Laplace`s equation? (b) calculate the total charge within the unit cube 0 < π‘₯, 𝑦, 𝑧 < 1π‘š.


1
Expert's answer
2021-01-20T07:24:29-0500

Answer

a) for laplace condition

"\nabla ^2V=0" ......... eq1

"\nabla ^2V=\frac{d^2V}{dx^2}+\frac{d^2V}{dy^2}+\frac{d^2V}{dz^2}"

"V=2x^2yz-y^3z"

"\frac{d^2V}{dx^2}=4yz"

"\frac{d^2V}{dy^2}=-6yz"

"\frac{d^2V}{dz^2}=0"

Putting all value in equation 1

"\nabla ^2V\ne0"

So this potential does not satisfy.

b) using poission equation

"\nabla ^2V=-\frac{\rho }{\epsilon}=-2yz"

Charge

"Q=\int\rho dv"

Integrate by putting charge density

"Q=\epsilon_0=8.85\times10^{-12}C"